OpenStudy (anonymous):

Rewrite each expression as an algebraic expression in terms of x. a) tan(arccosx) b) sin(2cos^-1(x)) c) tan(arccosx)

OpenStudy (zehanz):

a. Hint: use: tanx = sinx/cosx and sin²x+cos²x=1

OpenStudy (zehanz):

b. Do you mean:\[\cos^{-1} x = \frac{ 1 }{ \cos x }\]or\[\cos^{-1} x=\arccos x\]???